# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

### v3

@@ 5 Conjecture
Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type
@@
(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal
@@
In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 3.1 (page 3.1), and Remark 3.1 (page 3.1), Definition 2.4 (page 2.4), and Example 4.1 (page 4.1), with Diagram 4.3 (page 4.3). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
5.1 Conjecture: A Suitable Injection Does the Job of Characterisation
@@
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff
@@
5.2 Idea of Proof (Outline)
Fact 5.2
@@ 5.2 Idea of Proof (Outline)
We may also continue beyond hinaus: or then too.
@@
Fact 5.4
Adding new formulas inside a Yablo Triangle does not improve anything, it only destroys. It leads either to no contradiction, or to a contradiction of the simplest type, and thus to an escape situation, see Definition 2.4 (page 2.4).
There are no other new contradictory pairs of paths.
Proof
Consider a YC
Insert in ab, or and So, if then if then
Consider now insertions of xy, or xz, yz (analogously) and arbitrary combinations thereof.
xz: If then of then so we have a trivial contradiction and an escape possibility.
xy: xy as so If then if then we have again a trivial contradiction and an escape possibility.
yz: so If then if then we have again a trivial contradiction and an escape possibility.
Combinations:
For combinations of xy and yz, see (2.5) above.
xy and xz: the result is TRUE (both positive) or trivial contradiction with ESC.
xz and yz: This is analogous to xy and xz.
Let be such paths originating in containing some xy as in Fact 5.4 (page 5.4). Then the same considerations as above apply, so we have either no contradiction, or the trivial one with an escape possibility.
Similar results hold when using the constant TRUE: If has -TRUE in one branch, then will have TRUE in this branch, violating (C2). If has TRUE in one branch, then this branch is irrelevant, and we have no contradiction. So TRUE must not occur.
Extend the interpretation beyond
Let If has no successor, then set
If has successors etc., then set
If then has successors. Take the conjunction of the following: if if then
We have to show that still does what it should.
During construction, will be and after completion, And any will be TRUE or By the combination with AND, will not change.

### v4

@@ 4 Result
Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type It is important that these additions also cooperate with our truth value
@@
(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal nodes.)
@@
In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1), and Example 1.7 (page 1.7), with Diagram 1.2 (page 1.2). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
4.1 A Suitable Injection Does the Job of Characterisation
@@
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff the following conditions hold.
Condition 4.1
@@
4.2 Proof
Fact 4.2
@@ 4.2 Proof
We may also continue beyond : or then too.
@@
Fact 4.4
Consider (without the other elements etc. in
Now consider the full YS’, and attribute values to the nodes as follows, respecting the values already given in (1):
If is minimal, then
Otherwise:
If dann
arrow
Note that these arrows have no value, i.e. we have no distinction between and
If then
: arrow
: arrow
: arrow
Note that arrows inside have a value, these values have to be preserved. The other arrows have no value, and we treat them as positive.
Moreover, we did not distinguish between minimal and not minimal this does not matter. If we want to distinguish we have to differentiate the last conjunct further.
- (3)
Remark 4.5
Asssume for simplicity that (arrows may be negative, too), likewise for etc.
We have paths etc., assume for simplicity that the only new intermediate node is respectively).
Let is not in
We set arbitrary, Thus, so and so
Of course, instead of going through the additional path might also go through etc. This changes nothing. We have neutralised the path not in
Diagram 4.1
Diagram
Diagram 4.2
Diagram